OpenStudy (karategirl2002):

A train traveled 210 miles in 3.5 hours to its first stop. It then continued on traveling another 130 miles in 2.5 hours. What was the average speed of the train for the entire journey? A. 52 mi/h B. 56 mi/h C. 56.7 mi/h D. 60 mi/h

OpenStudy (anonymous):

Average speed=\[\frac{ total~ distance }{total~time }\]

OpenStudy (anonymous):

\[Average~ speed=\frac{ 210+130 }{3.5+2.5 }=?\]